# A Comprehensive Evaluation of Large Language Models on Mental Illnesses

### v1

@@ 3.3 Binary Disorder Classification (Task 1)
In the ZS binary disorder detection task, we evaluated the performance of various LLM s in identifying the presence or absence of depression, suicide risk, and stress in social media posts. We began by assessing their performance using the BIN-1 prompt template, followed by an analysis of the impact of modifications introduced in prompts BIN-2, BIN-3, and BIN-4.
@@ 3.3.1 Results of Prompt BIN-1
The Llama 2 70b model demonstrates notable performance, achieving 73.58% in DepSeverity, unexpectedly outperforming the Llama 3 70b (70.5%) and even surpassing GPT-4, which achieves 72.6%. The newer Llama 3 models exhibit greater consistency, with deviations of only 3-10% between the 8b and 70b variants, compared to 10-20% deviations in the Llama 2 family. However, the Llama 3.1 family of models was a significant disappointment, with the Llama 3.1 8b model performing considerably worse compared to the earlier version, except in two datasets (SAD & DEPTWEET), where the performance is generally comparable. A similar observation applies to the Llama 3.1 70b model, which is generally closer in performance to the earlier version but actually outperforms the Llama 3 70b model in the SAD and DEPTWEET datasets. Lastly, the Llama 3.1 405b model was a major disappointment, performing even worse than the Llama 3.1 70b model across all six datasets.
@@
Figure 4: Deviation of each model’s accuracy from the best model’s accuracy on Task 1 using Prompt BIN-1
@@ 3.3.2 Results of Prompts BIN-2/3/4
When examining the transitions from BIN-1 to BIN-3 and BIN-4, fewer models exhibited improved performance. Notably, models such as GPT-3.5-Turbo, Claude 3 Haiku, Llama 3 8b, Llama 2 70b, Llama 3.1 405b, Mistral NeMo, Phi-3-medium, and Qwen 2 72b showed declines ranging from 2% to over 20%, depending on the model and dataset. This trend indicates a general preference for the BIN-2 prompt, likely due to its clearer less structured approach.
@@
In terms of the overall best performing models, Gemma 7b as the top performer on the Dreaddit dataset with 76.70% accuracy (BIN-4), surpassing Mistral NeMo’s 74.50%. Other notable performances include Gemini 1.0 Pro, Mixtral 8x7b, and Mixtral 8x22b, which achieved accuracies of 72.20%, 69.50%, and 70.10% on BIN-4 and BIN-2, respectively. For the SDCNL dataset, Llama 3 70b matched GPT-4o’s 71.20% with BIN-2, while Gemma 7b reached 66.10% with prompt BIN-2. On the DepSeverity dataset, Mistral Medium achieved an accuracy of 75.10% (BIN-4), surpassing Llama 2 70b’s 73.58% (previous best on BIN-1). Several models, including Phi-3-small, Mixtral 8x7b, Mixtral 8x22b, Llama 3 70b, Gemma 2b, and Gemma 7b, approached or exceeded the 70% accuracy threshold, with Qwen 2 72b achieving the best performance at 76.60% (BIN-4).
@@ 3.4.1 Results of prompt SEV-4 on the ZS Severity Evaluation task
An analysis of Table 9 reveals that the performance of each model on the severity task.
The GPT family of models exhibits strong performance across multiple tasks, with GPT-4-Turbo leading the pack. This model achieves the highest BA of 39.6% in the DepSeverity task and an outstanding 59.7% in DEPTWEET, making it the top performer for these datasets. Its MAE scores are notably low, with the minimum being 0.455 in DEPTWEET. The average BA and MAE for GPT-4-Turbo stand at 41.4% and 1.047, respectively, highlighting its robustness and reliability. GPT-4 also performs admirably, with a notable BA of 40.9% in the RED SAM Test and 29.0% in SAD, securing the second-best score in the latter. Consistent MAE scores, such as 0.827 in DepSeverity and 0.751 in RED SAM, contribute to an average BA of 40.48% and MAE of 1.141, affirming its strong performance across various datasets. While GPT-4o doesn’t lead in any specific dataset, it demonstrates competitive results with a BA of 35.6% in DepSeverity and 50.1% in DEPTWEET, complemented by a low MAE of 0.825. The GPT-3.5-Turbo model delivers moderate performance, peaking at 43.2% BA in DEPTWEET and maintaining consistent MAE scores, including 0.888 in DepSeverity. With an average BA of 31.28% and MAE of 1.120, it aligns closely with GPT-4o, though it doesn’t reach the top-tier performance seen in the latest models. The GPT-4o mini, a recent addition to the series, falls slightly behind the larger GPT-4o model, achieving an average BA of 34.10% and MAE of 1.230, indicating a slight gap in performance despite its more compact architecture.
@@
Finally, the Qwen 2 72b model delivers moderate overall performance. While it doesn’t achieve standout results in any specific dataset, its performance is comparable to models like Llama 3 70b, Gemma 2 9b, or Phi-3-medium. It suggests a competent model that could benefit from further optimization to reach higher performance levels.
@@
Model DepSeverity DEPTWEET RED SAM Test SAD Average
@@ 3.4.2 Results of prompt SEV-4 on the FS Severity Evaluation task
Model DepSeverity DEPTWEET RED SAM Test SAD Average
@@ 3.5 Psychiatric Knowledge Assessment (task 3)
Llama 3.1 405b tops the performance chart, achieving an impressive accuracy of 91.2%, followed closely by Llama 3.1 70b with an accuracy of 89.50%. The smaller Llama 3.1 8b, however, lags behind with an accuracy of 68.0%. Despite this, it outperforms other models in its size category (under 14 billion parameters) such as Mistral 7b, Gemma 7b, and Phi-2, with a lead ranging from 8% to 20%. The older Llama 3 models all perform slightly worse compared to the more recently released models (e.g. Llama 3 70b actually beats GPT-4 even if only by 0.10%). Interestingly, it’s only in this task, which includes raw factual querying of the models, that the strength of the LLama 3.1 family and espcially the LLama 3.1 405b models stand out.
@@
Figure 5: Results of all models on Task 3.
@@
2022-11-282023-03-142023-07-182023-09-272023-11-142023-12-092023-12-122023-12-232024-02-212024-02-262024-03-142024-04-182024-05-132024-06-062024-06-202024-06-272024-07-022024-07-182024-07-232024-07-2430303030404040405050505060606060707070708080808090909090100100100100Claude 3 HaikuClaude 3 SonnetClaude 3.5 SonnetClaude 3 OpusGemma 2bGemma 7bGemma 2 9bGemma 2 27bGemini 1.0 ProGPT-4o miniGPT-3.5-TurboGPT-4oGPT-4-TurboGPT-4Llama 3 8bLlama 2 13bLlama 2 70bLlama 3 70bLlama 3.1 8bLlama 3.1 70bLlama 3.1 405bMistral 7bMistral NeMoMixtral 8x7bMixtral 8x22bMistral MediumMistral LargeMistral Large 2Phi-2Phi-3-miniPhi-3-smallPhi-3-mediumQwen 2 72bRelease DateAccuracyClaudeGemma/GeminiGPTLlamaMistralPhiQwen

### v2

@@ 4.2 Binary Disorder Classification (Task 1)
The results for the ZSbinary disorder detection task are presented below. We first examine the performance of models using the BIN-1 prompt, followed by an analysis of how modifications in prompts BIN-2, BIN-3, and BIN-4 affected model performance.
@@ 4.2.1 Results of Prompt BIN-1
The Llama 2 70b model demonstrates notable performance, achieving 73.58% in DepSeverity, unexpectedly outperforming the Llama 3 70b (70.5%) and even surpassing GPT-4, which achieves 72.6%. The newer Llama 3 models exhibit greater consistency, with deviations of only 3-10% between the 8b and 70b variants, compared to 10-20% deviations in the Llama 2 family. However, the Llama 3.1 family of models was a significant disappointment, with the Llama 3.1 8b model performing considerably worse compared to the earlier version, except in two datasets (SAD & DEPTWEET), where the performance is generally comparable. A similar observation applies to the Llama 3.1 70b model, which is generally closer in performance to the earlier version and actually outperforms the Llama 3 70b model in the SAD and DEPTWEET datasets. Lastly, the Llama 3.1 405b model was a major disappointment, performing even worse than the Llama 3.1 70b model across all six datasets.
@@
Figure 3: Average Deviation of Model Accuracy from the Best Performing Model’s Accuracy on each dataset of Task 1 (Dreaddit, SDCNL, DepSeverity, SAD, DEPTWEET, RED SAM) using Prompt BIN-1. Deviation is calculated for each dataset as the absolute difference in accuracy between a given model and the best performing model on that specific dataset. These deviations are then averaged across all datasets to produce the final values shown.
@@ 4.2.2 Results of Prompts BIN-2/3/4
When examining the transitions from BIN-1 to BIN-3 and BIN-4, fewer models exhibited improved performance. Notably, models such as GPT-3.5-Turbo, Claude 3 Haiku, Llama 3 8b, Llama 2 70b, Llama 3.1 405b, Mistral NeMo, Phi-3-medium, and Qwen 2 72b showed declines ranging from 2% to over 20%, depending on the model and dataset. This trend indicates a general preference for the BIN-2 prompt, likely due to its clearer, less structured approach.
@@
Regarding the overall best-performing models, Gemma 7b as the top performer on the Dreaddit dataset with 76.70% accuracy (BIN-4), surpassing Mistral NeMo’s 74.50%. Other notable performances include Gemini 1.0 Pro, Mixtral 8x7b, and Mixtral 8x22b, which achieved accuracies of 72.20%, 69.50%, and 70.10% on BIN-4 and BIN-2, respectively. For the SDCNL dataset, Llama 3 70b matched GPT-4o’s 71.20% with BIN-2, while Gemma 7b reached 66.10% with prompt BIN-2. On the DepSeverity dataset, Mistral Medium achieved an accuracy of 75.10% (BIN-4), surpassing Llama 2 70b’s 73.58% (previous best on BIN-1). Several models, including Phi-3-small, Mixtral 8x7b, Mixtral 8x22b, Llama 3 70b, Gemma 2b, and Gemma 7b, approached or exceeded the 70% accuracy threshold, with Qwen 2 72b achieving the best performance at 76.60% (BIN-4).
@@
4.2.3 Task 1 Invalid Response Rates
As shown in Figure 4, adherence to instruction formatting varied widely across the evaluated LLMs, with no clear correlation to parameter count. For example, the 8B Llama 3 8b (2.50%) and 3.8B Phi-3-mini (0.51%) matched or outperformed larger models like the 70B Llama 3 70b (0.37%) and GPT-4 (0.02%). Conversely, smaller architectures such as Phi-2 (2.7B, 26.98%) and Gemma 2b (2B, 33.53%) exhibited error rates far exceeding those of many large models, including Mistral Large (19.68%) and Claude 3 Opus (20.49%), which frequently appended explanations despite explicit formatting instructions.
The lowest invalid response rates occurred in newer models: Mistral NeMo (12B, 0.00%), Mistral Large 2 (0.00%), GPT-4o mini (0.00%), and GPT-4 (0.02%). Mid-sized architectures like Phi-3-medium (14B, 0.07%) and Gemma 2 27b (27B, 0.35%) achieved near-zero error rates comparable to much larger systems such as Llama 3 70b (70B, 0.37%) and GPT-3.5-Turbo (0.37%).
Meta’s Llama 3 series improves upon prior generations: Llama 3 70b (0.37%) and 8b (2.50%) outperform Llama 2 70b (0.43%) and 13b (6.37%). However, the newer Llama 3.1 70b (0.79%) and Llama 3.1 405b (9.00%) exhibit higher error rates than their base versions, likely owing to their new ethical guards (Section 7.1.2). Mistral’s Mistral 7b (7B, 12.29%) and Mistral Large (19.68%) are outperformed by successors Mistral NeMo and Mistral Large 2 (both 0.00%).
Claude models show substantial variability: Claude 3.5 Sonnet (0.80%) and Claude 3 Sonnet (1.15%) achieve lower rates than Claude 3 Haiku (15.90%) and Claude 3 Opus (20.49%). Mid-range performers include Gemini 1.0 Pro (1.37%), Qwen 2 72b (1.46%), GPT-4-Turbo (1.64%), Gemma 7b (1.88%), Phi-3-small (7B, 3.87%), Llama 3.1 8b (4.82%), and Mixtral 8x7b (6.08%).
The inverse relationship between model recency and error rates holds across most families, though exceptions exist. For example, Llama 3.1 405b (9.00%) and Llama 3.1 8b (4.82%) underperform relative to their release timing, diverging from the broader trend.
2022-11-282023-03-142023-07-182023-09-272023-11-142023-12-092023-12-122023-12-232024-02-212024-02-262024-03-142024-04-182024-05-132024-06-062024-06-202024-06-272024-07-022024-07-182024-07-232024-07-240.010.010.10.1111010100100Claude 3 HaikuClaude 3 SonnetClaude 3.5 SonnetClaude 3 OpusGemma 2bGemma 7bGemma 2 9bGemma 2 27bGemini 1.0 ProGPT-4o miniGPT-3.5-TurboGPT-4oGPT-4-TurboGPT-4Llama 3 8bLlama 2 13bLlama 2 70bLlama 3 70bLlama 3.1 8bLlama 3.1 70bLlama 3.1 405bMistral 7bMistral NeMoMixtral 8x7bMixtral 8x22bMistral MediumMistral LargeMistral Large 2Phi-2Phi-3-miniPhi-3-smallPhi-3-mediumQwen 2 72bRelease DateRate of Invalid Responses (Log Scale)
Figure 4: Invalid Response Statistics Sorted by model release date
@@ 4.3.1 Results of prompt SEV-4 on the ZSSeverity Evaluation task
An analysis of Table 10 reveals that the performance of each model on the severity task. The GPT family of models exhibits strong performance across multiple tasks, with GPT-4-Turbo leading the pack. This model achieves the highest BAof 39.6% in the DepSeverity task and an outstanding 59.7% in DEPTWEET, making it the top performer for these datasets. Its MAEscores are notably low, with the minimum being 0.455 in DEPTWEET. The average BAand MAEfor GPT-4-Turbo stand at 41.4% and 1.047, respectively, highlighting its robustness and reliability. GPT-4 also performs admirably, with a notable BAof 40.9% in the RED SAM Test and 29.0% in SAD, securing the second-best score in the latter. Consistent MAEscores, such as 0.827 in DepSeverity and 0.751 in RED SAM, contribute to an average BAof 40.48% and MAEof 1.141, affirming its strong performance across various datasets. While GPT-4o doesn’t lead in any specific dataset, it demonstrates competitive results with a BAof 35.6% in DepSeverity and 50.1% in DEPTWEET, complemented by a low MAEof 0.825. The GPT-3.5-Turbo model delivers moderate performance, peaking at 43.2% BAin DEPTWEET and maintaining consistent MAEscores, including 0.888 in DepSeverity. With an average BAof 31.28% and MAEof 1.120, it aligns closely with GPT-4o, though it doesn’t reach the top-tier performance seen in the latest models. The GPT-4o mini, a recent addition to the series, falls slightly behind the larger GPT-4o model, achieving an average BAof 34.10% and MAEof 1.230, indicating a slight gap in performance despite its more compact architecture.
@@
Finally, the Qwen 2 72b model delivers moderate overall performance. While it doesn’t achieve standout results in any specific dataset, its performance is comparable to models like Llama 3 70b, Gemma 2 9b, or Phi-3-medium. This indicates a competent model that could benefit from further optimization to reach higher performance levels.
@@
Model DepSeverity DEPTWEET RED SAM SAD Average
@@ 4.3.2 Results of prompt SEV-4 on the FSSeverity Evaluation task
Model DepSeverity DEPTWEET RED SAM SAD Average
@@ 4.4 Psychiatric Knowledge Assessment ( Task 3)
Llama 3.1 405b tops the performance chart, achieving an impressive accuracy of 91.2%, followed closely by Llama 3.1 70b with an accuracy of 89.50%. The smaller Llama 3.1 8b, however, lags behind with an accuracy of 68.0%. Despite this, it outperforms other models in its size category (under 14 billion parameters) such as Mistral 7b, Gemma 7b, and Phi-2, with a lead ranging from 8% to 20%. The older Llama 3 models all perform slightly worse compared to the more recently released models (e.g. Llama 3 70b actually beats GPT-4 even if only by 0.10%). Interestingly, it’s only in this task, which includes raw factual querying of the models, that the strength of the Llama 3.1 family and specifically the Llama 3.1 405b models stand out.
@@
Figure 5: Results of all models on Task 3 on the MedMCQA dataset
@@
2022-11-282023-03-142023-07-182023-09-272023-11-142023-12-092023-12-122023-12-232024-02-212024-02-262024-03-142024-04-182024-05-132024-06-062024-06-202024-06-272024-07-022024-07-182024-07-232024-07-243030404050506060707080809090100100Claude 3 HaikuClaude 3 SonnetClaude 3.5 SonnetClaude 3 OpusGemma 2bGemma 7bGemma 2 9bGemma 2 27bGemini 1.0 ProGPT-4o miniGPT-3.5-TurboGPT-4oGPT-4-TurboGPT-4Llama 3 8bLlama 2 13bLlama 2 70bLlama 3 70bLlama 3.1 8bLlama 3.1 70bLlama 3.1 405bMistral 7bMistral NeMoMixtral 8x7bMixtral 8x22bMistral MediumMistral LargeMistral Large 2Phi-2Phi-3-miniPhi-3-smallPhi-3-mediumQwen 2 72bRelease DateAccuracyClaudeGemma/GeminiGPTLlamaMistralPhiQwen
@@
4.5 Fine-tuning Exploration
This section presents and discusses the results of our initial fine-tuning explorations. Our primary goal here is to examine the in-domain and out-of-domain performance impact of fine-tuning GPT-4o-mini and Phi-3.5-mini on Disorder Severity Assessment.
To initially probe the benefits of fine-tuning, we conducted targeted experiments with both GPT-4o-mini and Phi-3.5-mini. Focusing on GPT-4o-mini, fine-tuning on a 1600-sample, balanced DEPTWEET subset yielded encouraging in-domain gains, achieving a notable 27.5 percentage point improvement in BAon DEPTWEET itself. Evaluating out-of-domain transfer, fine-tuned GPT-4o-mini also demonstrated positive performance shifts on DepSeverity in fine-tuned settings, with a 5.6% increase in BAaccompanied by a decrease of 0.028 in MAE. Similarly, in-domain fine-tuning on DEPTWEET resulted in a decreased MAEof 0.297 (decrease in MAEof 0.352). However, this positive trend was not universally observed; on the RED SAM dataset, fine-tuning on DEPTWEET led to a decrease in performance (-3.4% BAand an increase of +0.321 MAE).
A similar mixed trend emerged with Phi-3.5-mini, where we explored parameter-efficient fine-tuning employing LoRA in 4-bit quantization (Q4). Fine-tuning Phi-3.5-mini on DEPTWEET yielded substantial in-domain improvements (+45.5% BA& -0.595 MAEon DEPTWEET) and positive transfer to DepSeverity (+12.7% BA& -0.215 MAE) and Red Sam (+7.7% BA& -0.082 MAE), though a slight degradation was observed on SAD (-6% BA& +0.25 MAE). Interestingly, fine-tuning Phi-3.5-mini on the stress-focused SAD dataset, instead of DEPTWEET, resulted in a different performance profile, with more modest in-domain gains on SAD (+30.2% BA& -1.766 MAE), but also smaller, yet still positive transfer effects on DEPTWEET (+14.4% BA& -0.184 MAE), DepSeverity (+4.8% BA& -0.088 MAE), and RED SAM (+4% BA& -0.037 MAE).
Fine-tuning was performed using the LoRA method implemented in the Hugging Face Transformers library, with an Adam optimiser, a learning rate of 2e-5, a batch size of 16, and a maximum of 3 epochs. We used early stopping if validation performance, measured by balanced accuracy for classification tasks and MAE for severity estimation, did not improve for two consecutive epochs. Validation sets comprised 20% of the training data, stratified by outcome class or severity score. All experiments were run on NVIDIA A100 GPUs. We tracked both training cost (GPU-hours) and inference latency before and after fine-tuning.
4.6 Limitations
This study demonstrates the potential of applying LLMs to mental health tasks, but several limitations should be considered when interpreting the findings. First, the reliance on social media data introduces concerns regarding data quality, representativeness, and labeling subjectivity, all of which may bias results. In addition, practical constraints such as computational cost, context-length restrictions, APIfiltering, prompt sensitivity, and the varying degrees of “stubbornness” observed across different LLMs challenge both the scalability and generalizability of these methods. Beyond technical factors, the inherent subjectivity of psychiatric evaluation, coupled with ethical considerations related to data privacy, algorithmic bias, and model transparency, further complicates deployment in real-world settings. While the results highlight encouraging opportunities, they should be interpreted cautiously, with recognition of these challenges and the need for continued research. A more detailed discussion of these limitations and their implications for generalizability is presented in Section 7.1.
A further limitation lies in the absence of comprehensive statistical significance testing when comparing model performance across tasks and prompt conditions. Although prior work has underscored the importance of such testing to ensure robust interpretation—particularly in cases of modest performance differences—implementing it here was infeasible. This limitation stemmed from the large number of models evaluated ($\sim$35), restricted access to several APIs, and the prohibitive computational and financial costs of repeated experimental runs. Future work should prioritize integrating statistical significance testing frameworks as model access stabilizes and benchmarking resources expand, thereby enabling more rigorous and reliable comparative analysis.
5 Implications of the Study
This study provides a comprehensive look at how well different LLMs can perform on mental health tasks. The findings have important implications for using AIin psychiatry. In this section, we discuss the practical and ethical implications of deploying LLMs in this sensitive domain. THe implications can be summarized int he following points
  • Model Selection: There’s no "one-size-fits-all" model for mental health tasks. Performance varies significantly across datasets. Developers need to carefully choose models based on their specific needs and the type of data they’re working with. However, some models demonstrated more consistent performance across tasks than others. Notably, the OpenAI GPT family consistently ranked among the top performers in binary classification, severity estimation, and knowledge assessment. Similarly, the Mistral family, particularly Mistral NeMo and Mistral Medium, showed strong and consistent performance on certain datasets. As such, developers should test different models on data similar to theirs to find the best fit.
  • Closed vs Open-Source Models: While OpenAI’s models were the best overall models for many tasks, they remain closed-source, without the possibility of the stricter monitoring and transparency necessary for model deployment in this field. As such, open-source models offer a viable and scalable alternative for many mental health applications, especially considering the sensitive nature of the domain. Models like Llama 3 70b are comparable to closed-source models like GPT-4 in binary disorder detection, while being much smaller. Mistral NeMo, an even smaller model, is great for quick and efficient processing. They can be customized (i.e. via fine-tuning) and deployed more easily, especially when resources are limited. Furthermore, both open-source and closed-source models (where providers allow) can be fine-tuned on specific datasets. This can further improve performance and tailor the models to particular mental health tasks or populations.
  • Prompt Engineering is Key: Prompt engineering significantly improve model performance, especially for smaller models. The results presented in this study highlight the significant gains that can be achieved through careful prompt design and refinement. Rather than simply selecting a model and deploying it directly, developers should invest effort in systematically crafting and testing various prompts to identify those that result in the most accurate and reliable responses. As demonstrated by the substantial improvement of Gemma 7b, where prompt engineering elevated its accuracy on the Dreaddit dataset to 76.70% (an improvement of 22.58%) using the BIN-4 prompt, this process can unlock the full potential of models, even those with limited computational resources. Furthermore, our experiments, as shown in Table 13 in Section 6, suggest that incorporating a repetition line—a restatement of the core instruction—within the prompt can also enhance performance, potentially by reinforcing the desired behavior in the model.
  • FSLearning Improves Accuracy: FSlearning generally improves accuracy. This was evident across nearly all models and datasets tested, such as the significant performance boost observed in Phi-3-mini, achieving an average improvement of +8.18% in BAand a reduction of -0.488 in MAEacross the datasets tested. While the degree of improvement varied, the trend was clear: providing examples generally led to better results. However, it’s important to remember that FSlearning typically increases computational cost, often using 3-4x more tokens during inference (with three example posts provided, along with the one for which we require evaluation). As such, while FSlearning presents a powerful technique for enhancing model performance, developers must carefully weigh the benefits of improved accuracy against the increased computational demands, especially when deploying models in resource-constrained environments or when dealing with high-volume, real-time applications.
  • LLMs Can Estimate Severity, but with Limits: LLMs demonstrate a capability to estimate the severity of mental health conditions, but their accuracy is not yet at a level where they can be solely relied upon for critical decisions. While models like GPT-4-Turbo achieved a relatively low MAE of 0.455 on the DEPTWEET dataset, other models and datasets showed higher average errors. For instance, on the DepSeverity dataset, Mistral Medium had an MAE of 0.779, and on the SAD dataset, even the best-performing model, Claude 3 Sonnet, had an MAE of 1.404, which is relatively impressive on a 10 point severity scale. These results suggest that while LLMs can provide a general estimate of severity, often within one level of the actual value on a multi-level scale, they should be used in conjunction with, not in place of, human clinical judgment. The potential for error, as indicated by the MAE, should always be taken into account when interpreting model outputs.
  • Newer Models are Generally Better for Knowledge Tasks: Newer LLMs, such as the Llama 3.1 family, demonstrate a significantly improved understanding of psychiatric knowledge compared to their predecessors. In the psychiatric knowledge assessment task, Llama 3.1 405b achieved an impressive accuracy of 91.2% (exceeding GPT-4o), followed by Llama 3.1 70b at 89.50%. Even the smaller Llama 3.1 8b model outperformed many older, larger models. This suggests that newer models may have been trained on more comprehensive datasets or have architectures better suited for knowledge retention and retrieval. Therefore, for applications requiring an accurate and broad psychiatric knowledge base, such as question-answering, prioritizing the use of the most recent LLMs is likely to yield the best results. Developers should consider regularly updating their systems to incorporate advancements in newer model releases. Furthermore, techniques like Retrieval-Augmented Generation (RAG), while not explored in this study, could potentially further enhance performance by providing models with access to external, up-to-date knowledge sources.
  • Trade-offs Between Safety and Accuracy: Models with stricter ethical guidelines, like Llama 3.1 and Gemini, sometimes performed worse on mental health tasks. While these safeguards are intended to prevent harm and ensure responsible use, they can also limit the ability of LLMs to process and respond to sensitive information that may be crucial for accurate assessments. For example, the Llama 3.1 models exhibited a higher rate of invalid responses, potentially due to their heightened sensitivity to prompts related to mental health issues. As such, developers and researchers need to find a balance between making AIsafe and making it useful for mental health. Overly strict rules might limit the ability of LLMs to help with sensitive topics, while lax rules could lead to unintended negative consequences.
  • Bias is a Concern: While our experiments did not touch on this point, LLMs can inherit biases from their training data, potentially leading to unfair or inaccurate assessments, particularly for underrepresented groups. Although our results cannot directly determine whether a model’s incorrect answer stems from bias or other factors, developers must be aware of this risk and strive to mitigate potential biases.
  • Explainability is Important: It is often difficult to understand how LLMs arrive at their answers. This lack of transparency can be problematic in mental healthcare, where clinicians and patients need to understand the reasoning behind AI-generated assessments or recommendations. While true explainability remains a challenge, research into techniques like CoTprompting and attention visualization may offer insights into model decision-making.
  • Data Privacy Must be Protected: LLMs for mental health often use sensitive personal information. Thus, protecting data privacy is essential. Strict security measures, data anonymization techniques, and robust ethical guidelines are needed when collecting, storing, and using data for LLMtraining and deployment in mental healthcare.
In summary, this study demonstrates that while LLMs offer exciting possibilities for improving mental healthcare, realizing their full potential requires a balanced approach, careful model selection, thorough prompt engineering, and attention to ethical boundaries.